The first cost of installing new lighting along Maple Lane is $1,000,000. The new lighting will require maintenance of about $30,000 per year. This new illumination will improve the area and property values will be enhanced from $1,500,000 to $2,000,000 annually. The modified B/C ratio at a discount rate of 5% per year for a 15 year study period is closest to:

a. 0.045

b. 4.88

c. 0.48

d. 10.15
Answer #1

Initial Cost = $1,000,000

Annual Maintenance Cost = $30,000 per year

Discount rate = 5%

Study Period = 15 years

Savings is the enhance property values. Values increased from $1,500,000 to $2,000,000 annually.

Annual Savings = $2,000,000 - $1,500,000 = $500,000

Calculating Modified B/C ratio (Using PW Method)

Modified B/C ratio = PW of Benefits – PW of Maintenance Cost ÷ Initial Cost

Modified B/C ratio = $500,000 (P/A, 5%, 15) – $30,000 (P/A, 5%, 15) ÷ $1,000,000

Modified B/C ratio = $500,000 (10.37966) – $30,000 (10.37966) ÷ $1,000,000

Modified B/C ratio = $4,878,440 ÷ $1,000,000

Modified B/C ratio = 4.878 or 4.88

Answer – B) 4.88